Suppose that a price-searcher firm had consumers who were all identical to each other. The individual consumer's demand function is given by: qD= 40-4P. The firm decides to try a second-degree price discrimination scheme. The first 18 units will have a price of $5.50. After that, any units a consumer purchases will be only $2.75. The firm has a constant marginal cost of $1.00per unit.

Calculate the consumer surplus.
Answer #1

When P = 2.75, Q = 40 - 4*2.75 = 29

So total units from 18 to 29 are sold at Price = 2.75.

so total CS = CS1 + CS2

= [.5*(10-5.5)*18 ] + [ .5*(5.5-2.75)*(29-18)]

= 40.5 + 15.125

= 55.625